OpenStudy (anonymous):

A catapult launches a boulder with an upward velocity of 122 feet per second. The height of the boulder H in feet in after T seconds is given by the function H(t)= -16t^2 + 122t + 10. what is the boulders maximum height. how long does it take the boulder to reach it's maximum height

OpenStudy (mathstudent55):

At the moment of launch, t = 0, what height is the boulder at?

OpenStudy (anonymous):

0 I think

OpenStudy (mathstudent55):

Don't guess. Just plug in t = 0 in the equation. What do you get for h?

OpenStudy (anonymous):

i have no idea that's why im here asking for help

OpenStudy (mathstudent55):

You are given the equation \(h(t) = -16t^2 + 122t + 10\) which represents the height, h, of the boulder at time, t. Right?

OpenStudy (anonymous):

yes but when i add them together i get 116 and that is not right

OpenStudy (anonymous):

how or what do i need to do to get the right answer im not asking for the answer im asking how to get the answer

OpenStudy (mathstudent55):

You don't add anything together yet. Since we know the height as a function of time, we are going to find the height when the boulder is launched. At the very moment of the launch of the boulder, time is zero, t = 0. At t = 0, \(h(t) = -16t^2 + 122t + 10 \) \(h(0) = -16(0^2) + 122(0) + 10 \) \(h(0) = 10\) This means the boulder is at a height of 10 ft when it is launched.

OpenStudy (anonymous):

Would it be 242.56 feet after 3.81 seconds

OpenStudy (mathstudent55):

I don't know. I haven't solved the problem yet, and I have no answers to choose from.

OpenStudy (mathstudent55):